OpenStudy (anonymous):

Carol received 80, 84 and 75 on three Science tests. What did Carol receive on her fourth test if her average on all four tests was 82? Choose the correct equation and solution.

OpenStudy (anonymous):

(239 + x)/3 = 82; x = 89 (239 + x)/4 = 82; x = 89 (239 + x)/3 = 82; x = 85 (239 + x)/4 = 82; x = 85

OpenStudy (katherinesmith):

89

OpenStudy (anonymous):

the choices are right up there ^

OpenStudy (anonymous):

(239 +x)/4 = 82; x = 89